OpenStudy (anonymous):

PLEASE HELP ME!!!!! Find the discount. regular price = $8; discount rate = 75% $6 $0.60 $0.75 $2

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

I need help... please with a cherry on top?

OpenStudy (alexandervonhumboldt2):

multply the regular price by the discount rate to get the new price: then just subtract regular price - new price=discount

OpenStudy (alexandervonhumboldt2):

\[8-8*\frac{ 75 }{ 100 }=8*(1-0.75)=8*0.25=?\]

OpenStudy (anonymous):

kk thx!

OpenStudy (anonymous):

Can you telll me the answer?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

i can't give answers

OpenStudy (alexandervonhumboldt2):

just simpliy what i wrote

OpenStudy (anonymous):

plz...

OpenStudy (alexandervonhumboldt2):

never

OpenStudy (alexandervonhumboldt2):

it is not hard to 8*0.25=?

OpenStudy (anonymous):

the answer is $2... right?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (pyxieyt):

@Smart_Owl_Student I would like to tell you that we have an Anti-Cheating Policy! On OpenStudy you are not allowed to give the answers. You should guide the person through the problem. If we all just gave answers how would anyone learn. I do not want you to be in trouble for this is the future! (^o^)

OpenStudy (alexandervonhumboldt2):

the discount is 2

OpenStudy (anonymous):

Ok... thx

OpenStudy (alexandervonhumboldt2):

:)